OpenStudy (lovelyharmonics):

Find all solutions in the interval [0, 2π). 4 sin2x - 4 sin x + 1 = 0

OpenStudy (lovelyharmonics):

@Luigi0210

OpenStudy (luigi0210):

Do you think that's factorable?

OpenStudy (lovelyharmonics):

no.. becasue 1x4=4 and -1x-4=4 but -1-4=-5 and 1+4=5 so definitely no

OpenStudy (luigi0210):

Hmm, try the quadratic formula?

OpenStudy (lovelyharmonics):

+4sinx +-sqrt-4sinx^2-4(4sin2x)(1) all over 2(4sin2x)

OpenStudy (luigi0210):

Nope, just the constant, not the whole thing.

OpenStudy (lovelyharmonics):

what?

OpenStudy (lovelyharmonics):

like 4+-sqrt-4^2-4(4)(1) all over 2(4)?

OpenStudy (luigi0210):

Yes, and also you can make equations with http://prntscr.com/38jzw9

OpenStudy (luigi0210):

Makes it neater :)

OpenStudy (lovelyharmonics):

\[( 4+-\sqrt{0})/8 \]

OpenStudy (luigi0210):

Right! So that gives us \(x=.5\). Can you work backwards to get the original equation?

OpenStudy (lovelyharmonics):

(ノ_・。) .5 isnt an answer.....

OpenStudy (luigi0210):

Because we're not done yet silly >.>

OpenStudy (lovelyharmonics):

wait then what am i working backwards from? and i thought the sqrt of 0 was impossible

OpenStudy (luigi0210):

To get the equations \((2sinx-1)(2sinx-1)=0\)

OpenStudy (lovelyharmonics):

where did you get that out of (4+-sqrt0)/8

OpenStudy (luigi0210):

When you simplify \( \frac{4 \pm \sqrt{0}}{8}\) you get \(\frac{1}{2}\) when you simplify, right?

OpenStudy (lovelyharmonics):

i just went up a level for classifying a triangle (ノ◕ヮ◕)ノ*:・゚✧

OpenStudy (lovelyharmonics):

no, 4/8=1/2 and sqrt0/8= i have no idea

OpenStudy (luigi0210):

Congrats :) And you're thinking is rational. \(\sqrt{0}=0\) so \(\frac{\sqrt{0}}{8}=0\) So all you have is 1/2

OpenStudy (luigi0210):

But you can't separate them, you have to simplify the square root first.

OpenStudy (luigi0210):

Sorry, confusing you again? .-.

OpenStudy (lovelyharmonics):

yeah you can... as long as theres a plus or minus inbetween...

OpenStudy (luigi0210):

Sorry, used to doing it all as one.. but yea. \(\sqrt{0}=0\) And \(4 \pm 0\) will give you \(4\)

OpenStudy (lovelyharmonics):

so i was right (⊙ヮ⊙)

OpenStudy (luigi0210):

Yea you were xD I'm trying to think of what to do next..

OpenStudy (luigi0210):

Wait, nvm I know, I was thinking too hard! So we have the equation \((2sinx-1)^2=0\) Can you solve that?

OpenStudy (lovelyharmonics):

exactly! im about to chuck my computer at a wall.... (2sinx-1)(2sinx-1)= 4sinx^2-4sinx+1

OpenStudy (luigi0210):

Is your computer hatin on you? D: And no, SOLVE, not DISTRIBUTE >.<

OpenStudy (luigi0210):

So take the square root of both sides: \(2sinx-1=0\), now add one and divide. \(sinx=\frac{1}{2}\) Now, just find the values of sin that make it equal 1/2 and you're done :)

OpenStudy (lovelyharmonics):

i did /).(\ i used foil

OpenStudy (luigi0210):

Don't use FOIL, that'll undo all our work.. we have to set it equal to zero and solve for x.

OpenStudy (luigi0210):

Using your interval given.. \(\LARGE x=\frac{\pi}{3}\) and \(\LARGE x=\frac{2\pi}{3}\)